# Two Proofs of a Structural Theorem of Decreasing Minimization on Integrally Convex Sets

### v1

@@ Abstract
Keywords: discrete optimization, discrete convex analysis, integrally convex set, descreasing minimization, lexicographic optimization, Fenchel-type duality
@@ 1 Introduction
For $x=(2,5,2,1,3)$ and $y=(1,5,2,4,1)$, for example, we have $x{\downarrow}=(5,3,2,2,1)$ and $y{\downarrow}=(5,4,2,1,1)$. Since $x{\downarrow}_{1}=y{\downarrow}_{1}$ and $x{\downarrow}_{2}<y{\downarrow}_{2}$, we have $x<_{\rm dec}y$. For a given set $S$ of vectors, an element $x$ of $S$ is called decreasingly minimal (or dec-min) in $S$ if it is minimal in $S$ with respect to $\leq_{\rm dec}$, that is, if $x\leq_{\rm dec}y$ for all $y\in S$. In general, dec-min elements may not exist and are not uniquely determined. We denote the set of all dec-min elements of $S$ by ${\rm decmin}(S)$. The problem of finding a dec-min element of a given set $S$ is called the decreasing minimization problem on $S$.
@@
Let $S$ be an M-convex set. The set of dec-min elements of $S$ can be represented as
@@
The objective of this paper is to give two different proofs to Theorem 1.3. The first proof, given in Section 3, resorts to the Fenchel-type duality theorem for integrally convex functions established recently by the present authors [13]. Although this seems to be a nice application of a recent result in discrete convex analysis, it is also natural to ask whether such a machinery is inevitable to prove Theorem 1.3. For example, the original proofs [3, 5] of Theorems 1.1 and 1.2 for M-convex and M2-convex cases are based on standard tools in combinatorial optimization. In Section 4, we give the second elementary proof, which shows the existence of a box $[z,z^{\prime}]_{{\mathbb{Z}}}$ in the statement without relying on any tools other than the definition of integrally convex sets and the existence of a face $F$ on the basis of Farkas’ lemma.
@@ 3 Proof Based on Fenchel-type Duality
$\langle p,x\rangle=p_{1}x_{1}+p_{2}x_{2}+\cdots+p_{n}x_{n}$ $\qquad(p,x\in{\mathbb{R}}^{n}),$
@@
which is nonempty by the assumption that $S$ admits a dec-min element. On the other hand, the Fenchel-type duality (Theorem 3.1) show that
@@
Example 3.1 ([12, Example 13.15]).
The formula (3.2) is illustrated for a simple example. In particular, we identify the vector $p^{*}$ in the formula.
@@ 4.1 Bounding by a unit box
We first reduce our argument to the case where $x_{i}\neq y_{i}$ for all $i\in N$. Let $N^{\prime}=\{i\in N\mid x_{i}=y_{i}\}$. If $N^{\prime}\neq\emptyset$, let $\hat{N}=N\setminus N^{\prime}$ and consider $\hat{S}=\{z|_{\hat{N}}\mid z\in S,z|_{N^{\prime}}=x|_{N^{\prime}}\}$, $\hat{x}=x|_{\hat{N}}$, and $\hat{y}=y|_{\hat{N}}$. Then $\hat{S}$ is an integrally convex set, $\hat{x},\hat{y}\in{\rm decmin}(\hat{S})$, $\|\hat{x}-\hat{y}\|_{\infty}\geq 2$, and $\hat{x}_{i}\neq\hat{y}_{i}$ for all $i\in\hat{N}$. In the following we denote $(\hat{S},\hat{x},\hat{y})$ simply by $(S,x,y)$ and assume that
@@
$\{i\in N\mid x_{i}=\alpha_{1}-1,\ y_{i}=\alpha_{1}\}=\{i\in N\mid y_{i}=\alpha_{1}\},\quad|\{i\in N\mid y_{i}=\alpha_{1}\}|=|V_{1}|$ (4.13)
from the definition (4.3) of $p$ and $x{\downarrow}=y{\downarrow}$. Hence follows $|W_{1}|=2|V_{1}|$. ∎
@@
|$|\{i\in N^{\prime}\mid y_{i}=\alpha_{2}\}|=|N^{\prime}\cap V_{2}|$ (4.22)
from the definition (4.3) of $p$ and $x{\downarrow}=y{\downarrow}$. Hence follows $|W_{2}|=2|N^{\prime}\cap V_{2}|$. ∎
@@ 4.2 Determining a face
$p_{i}=\varphi(a_{i}+1)-\varphi(a_{i})\qquad(i=1,2,\ldots,n).$ (4.25)
Then we have
${\rm decmin}(S)\subseteq B^{\circ}\subseteq\arg\min(\Phi_{\rm rap}[-p]).$ (4.26)
@@
Hence follows ${\rm decmin}(S)=\arg\min\{\langle p,x\rangle\mid y\in S\cap B^{\circ}\}$. ∎
@@
This definition can also be expressed as $p^{*}=(p|_{N_{0}}+q,p|_{N_{1}})$, where $p|_{N_{0}}$ and $p|_{N_{1}}$ represent the restrictions of $p$ to $N_{0}$ and $N_{1}$, respectively. (We use similar notations also for other vectors.)
@@
Since $\min\{\langle p^{*},y\rangle\mid y\in S\cap B^{\prime}\}\geq\beta$ by Claims 4.14 and 4.15, we have $\langle p^{*},x^{k}\rangle\geq\beta$ for all $k$. Therefore, $\langle p^{*},z\rangle\geq\beta$, from which the claim follows. ∎
@@
    $\arg\min(\Phi_{\rm rap}[-p])=\{(1,1,0,1),(1,0,1,1)\}.$
    We thus have equality in the inclusion relation $B^{\circ}\subseteq\arg\min(\Phi_{\rm rap}[-p])$ in (4.26).

### v2

@@ Abstract
Keywords: discrete optimization, discrete convex analysis, integrally convex set, decreasing minimization, lexicographic optimization, Fenchel-type duality
@@ 1 Introduction
For $x=(2,5,2,1,3)$ and $y=(1,5,2,4,1)$, for example, we have $x{\downarrow}=(5,3,2,2,1)$ and $y{\downarrow}=(5,4,2,1,1)$. Since $x{\downarrow}_{1}=y{\downarrow}_{1}$ and $x{\downarrow}_{2}<y{\downarrow}_{2}$, we have $x<_{\rm dec}y$. For a given set $S$ of vectors, an element $x$ of $S$ is called decreasingly minimal (or dec-min) in $S$ if it is minimal in $S$ with respect to $\leq_{\rm dec}$, that is, if $x\leq_{\rm dec}y$ for all $y\in S$. In general, dec-min elements may not exist (e.g., $S={\mathbb{Z}}^{n}$) and are not uniquely determined (e.g., $S=\{(1,2),(2,1)\}$). We denote the set of all dec-min elements of $S$ by ${\rm decmin}(S)$. The problem of finding a dec-min element of a given set $S$ is called the decreasing minimization problem on $S$.
@@
An M-convex set $S$ has a dec-min element. The set of dec-min elements of $S$ can be represented as
@@
The objective of this paper is to give two different proofs to Theorem 1.3. The first proof, given in Section 3, resorts to the Fenchel-type duality theorem for integrally convex functions established recently by the present authors [13]. Although this seems to be a nice application of a recent result in discrete convex analysis, it is also natural to ask whether such a machinery is inevitable to prove Theorem 1.3. For example, the original proofs [3, 5] of Theorems 1.1 and 1.2 for M-convex and M2-convex cases are based on standard tools in combinatorial optimization. In Section 4, we give the second elementary proof, which shows the existence of a box $[z,z^{\prime}]_{{\mathbb{Z}}}$ in the statement without relying on any tools and the existence of a face $F$ on the basis of Farkas’ lemma.
@@ 3 Proof Based on Fenchel-type Duality
$\langle p,x\rangle=p_{1}x_{1}+p_{2}x_{2}+\cdots+p_{n}x_{n}\qquad(p\in{\mathbb{R}}^{n},x\in{\mathbb{Z}}^{n}),$
$\Phi[-p](x)=\Phi(x)-\langle p,x\rangle=\Phi(x)-\sum_{i=1}^{n}p_{i}x_{i}\qquad(p\in{\mathbb{R}}^{n},x\in{\mathbb{Z}}^{n}),$
$\arg\min\Phi[-p]=\{x\in{\mathbb{Z}}^{n}\mid\Phi[-p](x)\leq\Phi[-p](y)\mbox{ for all $y\in{\mathbb{Z}}^{n}$}\}\qquad(p\in{\mathbb{R}}^{n}),$
@@
which is nonempty by the assumption that $S$ admits a dec-min element. On the other hand, the Fenchel-type duality (Theorem 3.1) shows that
@@
Example 3.1.
The formula (3.2) is illustrated for a simple example [12, Example 13.15]. In particular, we identify the vector $p^{*}$ in the formula.
@@ 4.1 Bounding by a unit box
We first reduce our argument to the case where $x_{i}\neq y_{i}$ for all $i\in N$. Let $N^{\prime}=\{i\in N\mid x_{i}=y_{i}\}$. If $N^{\prime}\neq\emptyset$, let $\hat{N}=N\setminus N^{\prime}$ and consider
$\hat{S}=\{z|_{\hat{N}}\mid z\in S,z|_{N^{\prime}}=x|_{N^{\prime}}\},\qquad\hat{x}=x|_{\hat{N}},\qquad\hat{y}=y|_{\hat{N}},$
where, for any vector $z$ on $N$ and any subset $U$ of $N$, $z|_{U}$ denotes the restriction of $z$ to $U$. Then $\hat{S}$ is an integrally convex set, $\hat{x},\hat{y}\in{\rm decmin}(\hat{S})$, $\|\hat{x}-\hat{y}\|_{\infty}\geq 2$, and $\hat{x}_{i}\neq\hat{y}_{i}$ for all $i\in\hat{N}$. In the following we denote $(\hat{S},\hat{x},\hat{y})$ simply by $(S,x,y)$ and assume that
@@
$\{i\in N\mid x_{i}=\alpha_{1}-1,\ y_{i}=\alpha_{1}\}=\{i\in N\mid y_{i}=\alpha_{1}\},\quad|\{i\in N\mid y_{i}=\alpha_{1}\}|=|V_{1}|,$ (4.13)
since $\{i\in N\mid x_{i}\leq\alpha_{1}-2,\ y_{i}=\alpha_{1}\}=\emptyset$ from the definition (4.3) of $p$ and the assumption (A1). Hence follows $|W_{1}|=2|V_{1}|$. ∎
@@
|$|\{i\in N^{\prime}\mid y_{i}=\alpha_{2}\}|=|N^{\prime}\cap V_{2}|,$ (4.22)
since $\{i\in N^{\prime}\mid x_{i}\leq\alpha_{2}-2,\ y_{i}=\alpha_{2}\}=\emptyset$ from the definition (4.3) of $p$ and the assumption (A1). Hence follows $|W_{2}|=2|N^{\prime}\cap V_{2}|$. ∎
@@ 4.2 Determining a face
$p_{i}=\varphi(a_{i}+1)-\varphi(a_{i})\qquad(i=1,2,\ldots,n),$ (4.25)
for which $\arg\min(\varphi[-p_{i}])=\{a_{i},a_{i}+1\}$ for $i=1,2,\ldots,n$. Then we have
${\rm decmin}(S)\subseteq S\cap B^{\circ}\subseteq S\cap\arg\min(\Phi_{\rm rap}[-p]).$ (4.26)
@@
Hence follows ${\rm decmin}(S)=\arg\min\{\langle p,y\rangle\mid y\in S\cap B^{\circ}\}$. ∎
@@
This definition can also be expressed as $p^{*}=(p|_{N_{0}}+q,p|_{N_{1}})$, where $p|_{N_{0}}$ and $p|_{N_{1}}$ represent the restrictions of $p$ to $N_{0}$ and $N_{1}$, respectively.
@@
Since $\min\{\langle p^{*},y\rangle\mid y\in S\cap B^{\prime}\}\geq\beta$ by Claim 4.15, we have $\langle p^{*},x^{k}\rangle\geq\beta$ for all $k$. Thus we obtain $\langle p^{*},z\rangle\geq\beta$, while $\langle p^{*},z\rangle=\langle p^{*},(1-\varepsilon)x^{\rm c}+\varepsilon y\rangle=(1-\varepsilon)\beta+\varepsilon\langle p^{*},y\rangle$. Therefore $\langle p^{*},y\rangle\geq\beta$, from which the claim follows. ∎
@@
    $S\cap\arg\min(\Phi_{\rm rap}[-p])=\{(1,1,0,1),(1,0,1,1)\}.$
    We thus have equality in the inclusion relation $S\cap B^{\circ}\subseteq S\cap\arg\min(\Phi_{\rm rap}[-p])$ in (4.26).